I am having problem,s with this problem
Solve the following equation by factoring or by any of the techniques.
4x^2-8x+3=5
Thank a million for the help

Answer by stanbon(75887)

Solve the following equation by factoring or by any of the techniques.
4x^2-8x+3=5
Rearrange:
4x^2-8x-2=0
Divide thru by 2
2x^2-4x-1=0
Use the quadratic formula to get:
x = [4+-sqrt(16-4*2*-1)]/4
x = [4+-sqrt(24)]/4
x = [4+-2sqrt(6)]/4
x = 1+(1/2)sqrt(6) or x = 1 -(1/2)sqrt(6)
================
Cheers,
Stan H.